I need to insert parentheses in the equation to make it true:
30 divided by 2 + 4 x 2 + 7 = 10

and
30 divided by 2 + 4 x 2 + 7 = 51

Answer by Moni(1)

30/(2+4x2)+7=30/(2+8)+7=30/10+7=3+7=10
30/2+4x(2+7)=15+4x9=15+36=51